OpenStudy (anonymous):

If the sum of the interior angles of a polygon is 720 degrees how many sides does the polygon have?

OpenStudy (anonymous):

What comes next? @PINKJELLYFISH

OpenStudy (johnweldon1993):

Sum = 180 ( n - 2) ....where 'n' = number of sides we have the sum and want 'n' so 720 = 180 (n - 2) Divide both sides by 180 4 = n - 2 what does 'n' equal...?

OpenStudy (anonymous):